OpenStudy (anonymous):

A force of 2i + 7j N acts on a body of mass 5kg for 10 seconds. The body was initially moving with constant velocity of i – 2j m/s. Find the velocity of the body in m/s, in vector form.

OpenStudy (anonymous):

use newtons second law force = mass * acceleration

OpenStudy (anonymous):

i did not get it pls

OpenStudy (anonymous):

2i + 7j = 5 * a acceleration a = 2i/5 + 7j/5 noe use equations of constant acceleration velocity = v , u = i - 2j, a = 2i/5 + 7j/5, t = 10 v = u + at v = i - 2j + 10(2i/5 + 7j/5) = i - 2j + 4i + 14j = 5i + 12j